Toselli Animation plans to offer its employees a salary enhancement package that has revenue sharing as its main component. Specifically, the company will set aside 1% of total sales revenue for year-end bonuses. The sales are expected to be $5 million the first year, $5.5 million the second year, and amounts increasing by 10% each year for the next 5 years. At an interest rate of 7% per year, what is the equivalent annual worth in years 1 through 5 of the bonus package?

The equivalent annual worth of the bonus package is $
Answer #1

Present worth = 50000[1-(1+0.1/1+0.07)^5/0.07-0.1]

= 50000*1-1.1481/-0.03

= 50000*4.9367

= 246835

Equivalent annual worth = 246835*(A/P,i,n)

= 246835*(A/P,7%,5)

= 246835*(0.2439)

= 60093.30